# Thinking Before Running! Efficient Code Generation with Thorough Exploration and Optimal Refinement

### v1

@@ 4.2 Model Training and Inference
Generating high-probability text does not necessarily align well with human preferences, which may lead to biases in various tasks Perez et al. (2022). Reinforced Self-Training(ReST) mitigates this issue effectively by collecting offline data aligned with human preferences for online training Gulcehre et al. (2023). We train the base model $\mathcal{M}$ with the success trajectory dataset $D_{M_{0}}\cup R_{M_{0}}$:
[…]
In the inference process, two distinct configurations are used. The first configuration utilizes the fine-tuned base model, where the input $x$ is fed directly into the model to generate text with a temperature setting of $t=0$. This setup is designed to assess whether the data collected through ThinkCoder enhances the quality of the generated output. The second configuration regards the fine-tuned base model as the exploration agent within the ThinkCoder framework. This approach generates and refines solutions using $k$ explorations and $n$ refinements, with various temperature settings $t$. It aims to investigate the framework’s ability to improve the generated solutions through the success trajectories and the ThinkCoder framework. Comprehensive experimental results are presented in the evaluation section.
@@ 5.1 Dataset
We utilized the MBPP Austin et al. (2021) and HumanEval Chen et al. (2021) datasets to evaluate the effectiveness of our approach. To ensure a comprehensive assessment, we tested their enhanced versions—MBPP-ET and HumanEval-ET Dong et al. (2023)—which include 80 and 35 times more test cases than the original datasets, respectively. Regarding the test data, the original problem set sizes for MBPP and HumanEval are 257 and 164, while their extended versions contain 378 and 164 problems, respectively. For training, we started with 374 MBPP training examples and 90 validation examples, combining them as the initial dataset. Using temperature-based and reflection-based generation methods, we expanded this set to over 2,000 examples for ReST training.
@@ 5.2 Baselines
In terms of LLM base model selection for the application of ThinkCoder, we considered a range of advanced instruction-tuned models, including the general-purpose models LLama2-7B-Chat, Kimi [^1] https://platform.moonshot.cn, GPT-4-Turbo(gpt-4-1106-preview) and GPT-4o, as well as the code-specific model CodeQwen1.5-7B-Chat. We conducted a comprehensive evaluation of ThinkCoder’s Pass@1 performance on MBPP and HumanEval across LLMs of different parameter sizes. For the test-time optimization process, we used the base model LLama2-7B to evaluate performance changes on the MBPP dataset. Comparisons were made between results obtained before and after fine-tuning, using either the original training data or the success trajectories, and incorporating the ThinkCoder framework or excluding it. For the composite model, we compared it with the current State-of-the-Art (SOTA) models, Self-Collaboration Dong et al. (2024), MapCoder Islam et al. (2024), and AgentCoder Huang et al. (2023b).
@@
5.3 Implementation details
We set different parameter combinations for different experimental settings. To validate the improvement of the base LLMs, we set the exploration budget $k=5$, the temperature received by the exploration agent $t=0.5$, the number of optimal refinements $n=5$, and the budget control threshold $\theta=0.8$. The exploration agent generates $m=3$ test cases per code. To compare ThinkCoder with state-of-the-art (SOTA) results, we set the hyperparameters $k$, $t$, $n$, $\theta$, and $m$ to 20, 1.0, 2, 1.0, and 3, respectively. In the data collection phase, the exploration agent $\mathcal{M}_{0}$ is specified as CodeQwen1.5-7B-Chat because it is free and performs well. During ReST training, we set the learning rate to $2\times 10^{-5}$ and train for 2 epochs. We employ Low-Rank Adaptation (LoRA) Hu et al. (2021) to train LLMs.
6 Evaluation
Table 1 illustrates the performance changes of various base LLMs on the MBPP and HumanEval datasets as the Optimal Refinement iteration $n$ increases. Table 3 compares the performance of ThinkCoder and SOTA models, using GPT-4-Turbo as the base LLM. Table 2 presents the performance changes of LLama2-7B before and after fine-tuning with trajectory data collected by ThinkCoder. Figure 3 shows the impact of changes in the hyperparameter temperature $t$, exploration budget $k$, and refinement iterations $n$ on the Pass@1 and Pass@k metric. Figure 4 depicts the trends in the accuracy of generated test cases and generated codes as the optimal refinement iterations in the ThinkCoder increase. Table 4 compares ThinkCoder’s computation load with MapCoder at similar settings.
[…]
Self-Collaboration 90.2 70.7 78.9 62.1 75.5
MapCoder 83.1 57.7 93.9 82.9 79.4
AgentCoder 91.4 91.4 89.6 76.2 87.2
ThinkCoder 90.5 80.5 93.9 86.0 87.7
Table 3: End-to-end results of ThinkCoder and state-of-the-art (SOTA) baselines.
6.1 Performance on Code Generation
Models Exploration k HumanEval MBPP
Pass@1 Refinement n Tokens (k) Pass@1 Refinement n Tokens (k)
MapCoder 5 93.9 5 10.41 83.1 3 4.96
ThinkCoder 5 90.2 1 1.92 89.5 1 1.44
@@ 6.4 Hyperparameters study
We conducted hyperparameter selection experiments on the MBPP dataset. As shown in Figure 3(a), models exhibit optimal performance at different temperature settings. LLama2-7B-Chat, CodeQwen1.5-7B-Chat and Kimi improve in Pass@1 as temperature increases, while GPT-4-Turbo and GPT-4o perform better at lower temperatures. As depicted in Figure 3(b), the Pass@k performance of various base LLMs improves with an increasing exploration budget $k$, highlighting the models’ ability to find optimal solutions. However, the performance eventually levels off, suggesting that the Pass@k performance of these models will eventually stabilize. As shown in Figure 3(c), Pass@1 performance of base LLMs improves with increasing refinement iterations $n$, following a trend similar to the budget curve during exploration. However, exploration consistently achieves greater performance gains, underscoring the importance of optimizing the LLM’s answer selection ability for better code quality. For the hyperparameter study of $\theta$ and $m$, please refer to Appendix A.1.
@@ 6.5 Cost Analysis
In Figure 4, we compare the computational overhead between MapCoder and ThinkCoder. On the HumanEval and MBPP datasets, a single refinement with ThinkCoder outperforms the result of five iterations of refinement with MapCoder, while using fewer tokens. This is because MapCoder utilizes four agents that are all powered by LLMs, whereas our framework uses a single LLM to handle both code and test generation, while debugging is directly based on feedback from the Python environment, significantly reducing token usage.
@@ 7 Conclusion
In this paper, we introduce ThinkCoder, a framework designed to improve the code generation capabilities of LLMs. The proposed approach incorporates an extensive exploration to generate diverse solutions, followed by an optimal refinement stage that integrates the solutions, tests, and feedback as reflections to enhance performance. To minimize computational costs during exploration, we capture successful trajectories through ThinkCoder, which are then leveraged to fine-tune the LLM. This process not only boosts the model’s performance in code generation but also reduces the number of refinement iterations needed to achieve optimal solutions. We evaluate our framework across LLMs of varying sizes and demonstrate its superior performance over current composite models.
@@
In our current approach to managing the testing pool, we primarily focus on selective expansion but lack effective strategies for cleaning noise from the test set. Moreover, the diversity of the test set heavily depends on the generation capabilities of LLMs. Our work highlights the importance of tests and identifies the gap between generated and annotated cases, so we plan to address the enhancement of the testing pool in future work. Furthermore, while ReST training has demonstrated excellent results in enhancing the code generation capabilities of foundational LLMs, validation in ThinkCoder reveals rapid convergence. This indicates that fine-tuning teaches the LLM to find good answers, but the improvement in exploring the overall solution space remains limited. To address this, we look forward to introducing on-policy training methods in future work, combining data collection with fine-tuning to enable the model to evolve continuously during training, generating more diverse answers and further boosting the LLM’s online exploration capabilities.
@@ Appendix A Appendix
A.1 Exploration Budget Control
@@
A.2 Exploration for Tests
@@
Table 5: The variation in the Pass@1 metric and the execution agent’s overload is influenced by changes in the number of test cases $m$ during each exploration.
We randomly selected 50 tasks from the MBPP test data to observe the performance improvement and code execution overhead as the number of test cases $m$ increases during each exploration. Table 5 illustrates this trend. The base model we used is Kimi. We observed that increasing $m$ leads to improved Pass@1 performance and results in longer execution times due to the additional effort required for self-verification. To balance time efficiency and performance, we set $m$ for all experiments.
A.3 Prompts for Exploration Agent
@@
A.4 Prompts for Execution Agent
Figure 7: ThinkCoder execution agent prompt example.
@@ A.4 Prompts for Execution Agent
A.5 Example of Thorough Exploration
@@
A.6 Example of Optimal Refinement
@@
A.7 Example of Testing Pool

### v2

@@ 4.2 Model Training and Inference
Generating high-probability text does not necessarily align well with human preferences, which may lead to biases in various tasks Becker et al. (2024). Reinforced Self-Training(ReST) mitigates this issue effectively by collecting offline data aligned with human preferences for online training Gulcehre et al. (2023). We train the base model $\mathcal{M}$ with the success trajectory dataset $D_{M_{0}}\cup R_{M_{0}}$:
$\textstyle\mathcal{L}_{MLE}(\theta_{\mathcal{M}})=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{\mathcal{M}_{0}}\cup\mathcal{R}_{\mathcal{M}_{0}}}\log p_{\theta_{\mathcal{M}}}(y\mid x)$
In inference, we compare the performance with baselines at two configurations. The first uses a fine-tuned base model, feeding input $x$ directly to generate text at temperature $t=0$. This setup tests if ThinkCoder data improves output quality. The second configuration uses the fine-tuned base model as the exploration agent in ThinkCoder. It generates and refines solutions with $k$ explorations and $n$ refinements, using different temperature settings $t$. This aims to assess the framework’s ability to improve solutions through successful trajectories and the ThinkCoder framework.
@@ 5.1 Dataset
We utilized the MBPP Austin et al. (2021) and HumanEval Chen et al. (2021) datasets to evaluate the effectiveness of our approach. To ensure a comprehensive assessment, we tested their enhanced versions—MBPP-ET and HumanEval-ET Dong et al. (2023)—which include 80 and 35 times more test cases than the original datasets, respectively. Regarding the test data, the original problem set sizes for MBPP and HumanEval are 257 and 164, while their extended versions contain 378 and 164 problems, respectively. For training, we started with 374 MBPP training examples and 90 validation examples, combining them as the initial dataset. Using temperature-based and reflection-based generation methods, we expanded this set to over 2,000 examples for ReST training. We also evaluate our framework on two major benchmarks designed for harder problems: LiveCodeBench and CodeContests. LiveCodeBench consists of the 167 most recent samples collected between October 2024 and January 2025, while CodeContests includes 165 problems.
@@ 5.2 Baselines
In terms of LLM base model selection for the application of ThinkCoder, we considered a range of advanced instruction-tuned models, including the general-purpose models LLama2-7B-Chat, Kimi [^2] https://platform.moonshot.cn, GPT-4-Turbo(gpt-4-1106-preview) and GPT-4o(gpt-4o-2024-05-13), as well as the code-specific model CodeQwen1.5-7B-Chat. We conducted a comprehensive evaluation of ThinkCoder’s Pass@1 performance on MBPP and HumanEval across LLMs of different parameter sizes. For the test-time optimization process, we used the base model LLama2-7B to evaluate performance changes on the MBPP dataset. We compared results obtained before and after fine-tuning, using either the original training data or the success trajectories and incorporating or excluding the ThinkCoder framework. For the composite model, we compared it with the current SOTA models, MapCoder Islam et al. (2024), AgentCoder Huang et al. (2023b), LDB Zhong et al. (2024), and LPW Lei et al. (2024).
[…]
GPT-4-Turbo 0 33.4 16.1
1 43.7(30.8%) 26.1(62.1%)
2 46.1(38.0%) 27.3(69.6%)
5 44.9(34.4%) 28.5(77.0%)
GPT-4o 0 22.5 21.0
1 35.3(56.9%) 34.1(62.4%)
2 38.3(70.2%) 38.9(85.2%)
5 41.3(83.6%) 40.1(91.0%)
Table 4: ThinkCoder performance on LiveCodeBench and CodeContests. $n$ represents the self-refinement budget, with the self-exploration budget fixed at $k=5$.
5.3 Implementation details
We set different parameter combinations for different experimental settings. To validate the improvement of the base LLMs, we set the exploration budget $k=5$, the temperature received by the exploration agent $t=0.5$, the number of optimal refinements $n=5$, and the budget control threshold $\theta=0.8$. The exploration agent generates $m=3$ test cases per code. To compare ThinkCoder with SOTA results, we set the hyperparameters $k$, $t$, $n$, $\theta$, and $m$ to 20, 1.0, 2, 1.0, and 3, respectively. In the data collection phase, the exploration agent $\mathcal{M}_{0}$ is specified as CodeQwen1.5-7B-Chat because it is free and performs well. During ReST training, we set the learning rate to $2\times 10^{-5}$ and train for 2 epochs. We employ Low-Rank Adaptation (LoRA) Hu et al. (2021) to train LLMs.
6 Evaluation
Table 1 illustrates the performance changes of various base LLMs on the MBPP and HumanEval datasets as the Optimal Refinement iteration $n$ increases. Table 2 and Table 3 compare the performance of ThinkCoder and SOTA models, using GPT-4-Turbo and GPT-4o as the base LLMs. Table 4 shows great potential on harder benchmarks. Table 5 compares ThinkCoder’s computation load with MapCoder at similar settings. Table 6 presents the performance changes of LLama2-7B before and after fine-tuning with trajectory data collected by ThinkCoder. Table 7 illustrates the ablation study of self-exploration and self-refinement. Figure 3 shows the impact of changes in the hyperparameter temperature $t$, exploration budget $k$, and refinement iterations $n$ on the Pass@1 and Pass@k metric. Figure 4 depicts the trends in the accuracy of generated test cases and codes as the optimal refinement iterations in the ThinkCoder increase.
6.1 Performance on Code Generation
Models HumanEval Mbpp CodeContest
n k Pass@1$\uparrow$ Tokens (k)$\downarrow$ n k Pass@1$\uparrow$ Tokens (k)$\downarrow$ n k Pass@1$\uparrow$ Tokens (k)$\downarrow$
MapCoder 5 5 93.9 21.3 3 3 83.1 5.58 3 3 28.5 18.3
ThinkCoder(Ours) 20 1 93.9 0.26 5 1 89.5 0.24 5 5 28.5 2.94
6.2 Ablation Study of Self-exploration and Self-refinement
Figures 3(b) and 3(c) highlight the performance gains from both self-exploration and self-refinement. The gain from self-exploration is greater than that from self-refinement. As illustrated in Table 7, self-refinement significantly enhances efficiency when paired with self-exploration. However, increasing the exploration budget excessively (e.g., $k=20$) leads to inefficiency. A moderate increase in the exploration budget accelerates the optimization of the self-refinement process.
@@ 6.3 Performance with ReST Training
Self-exploration Budget k Self-refinement Budget n
1 2 3 5 10 20
1 91.7 90.0 91.7 93.3 93.3 95.0
2 93.3 95.0 95.0 96.7 96.7 96.7
5 93.3 93.3 95.0 95.0 96.7 96.7
10 96.7 96.7 96.7 96.7 96.7 96.7
20 96.7 96.7 96.7 96.7 96.7 96.7
Table 7: The performance trends on the randomly 60 Mbpp samples on GPT-4 with varying compositions of self-exploration and self-refinement budgets.
@@ 6.5 Hyperparameters Study
We conducted hyperparameter selection experiments on the MBPP dataset. As shown in Figure 3(a), models exhibit optimal performance at different temperature settings. LLama2-7B-Chat, CodeQwen1.5-7B-Chat, and Kimi improve in Pass@1 as temperature increases, while GPT-4-Turbo and GPT-4o perform better at lower temperatures. Figure 3(b) indicates that Pass@k performance of base LLMs improves with increasing exploration budget $k$ but eventually stabilizes. Figure 3(c) shows that Pass@1 performance improves with more refinement iterations $n$, but exploration gains are more significant. For details on hyperparameters $\theta$ and $m$, see Appendix A.2 and A.3.
@@ 6.6 Cost Analysis
In Table 5, we compare the computational overhead between MapCoder and ThinkCoder. Compablue to MapCoder, which uses four LLM-based agents executed serially, ThinkCoder employs only one LLM-based agent and one non-LLM-based CodeVerifier, resulting in lower time overhead. Therefore, we use token overload as a measure of cost consumption. It reveals that ThinkCoder demonstrates more efficient resource utilization compared to MapCoder with a similar pass@1.
@@ 7 Conclusion
In this paper, we introduce ThinkCoder, a framework designed to improve the code generation efficiency of LLMs. The proposed approach incorporates LLM’s exploration with a non-LLM-based CodeVerifier that greatly reduces the token usage. To further minimize computational costs during exploration, we capture successful trajectories through ThinkeCoder to fine-tune the LLM. This process enhances code generation performance while reducing refinement iterations for optimal solutions. We evaluate our framework across LLMs of varying sizes and demonstrate its superior performance over current composite models.
Acknowledgments
This work is also supported by the Public Computing Cloud, Renmin University of China and by fund for building worldclass universities (disciplines) of Renmin University of China.
@@ Acknowledgments
In our current approach to managing the testing pool, we primarily focus on selective expansion but lack effective strategies for cleaning noise from the test set. Moreover, test set diversity relies heavily on LLMs’ generation capabilities. Our work highlights the importance of tests and the gap between generated and annotated cases. We plan to enhance the testing pool in future work. Furthermore, while ReST training has demonstrated excellent results in enhancing the code generation capabilities of foundational LLMs, validation in ThinkCoder reveals rapid convergence. This indicates that fine-tuning teaches the LLM to find good answers, but the improvement in exploring the overall solution space remains limited. To address this, we look forward to introducing on-policy training methods in future work, combining data collection with fine-tuning to continuously evolve the model during training. This will generate more diverse answers and enhance the LLM’s online exploration capabilities.
@@ Appendix A Appendix
A.1 Algorithm of ThinkCoder
Algorithm 1 shows the pseudo-code of ThinkCoder.
A.2 Exploration Budget Control
@@
Algorithm 1 ThinkCoder
1:Initialize the testing pool TP=[], set the globally max pass rate $g_{p}=0$ and best solution $g_{s}=\text{None}$
2:Set the hyperparameters $t=0.5,k=5,n=5,m=3,\theta=0.8$
3:Set the instruction prompt for code generation task and instruction for task $p$ as $I=[p,\text{prompt}]$
4:for each $iter\in n$ do
5:     Initialize the locally max pass rate $l_{p}=0$ and best solution ${l_{s}}={\text{None}}$ for iteration $iter$
6:     Randomly select $m$ successful test cases from TP as tests
7:     Select a new type of error feedback for test $f_{t}$ from TP as $f$
8:     if $f_{t}$ and $f$ is not None then
9:         Update the instruction $I=[p,g_{s},f_{t},f,\text{tests},\text{prompt}]$
10:     end if
11:     Get code solutions $\text{G}^{\prime}$ and test cases $\text{TP}^{\prime}$ from exploration agent for the instruction $I$ within $k$ iterations
12:     Deduplicate testing pool TP’ by AST
13:     for each $i\in|\text{G}^{\prime}|$ do
14:         Get the pass rate $r_{g_{i}}$ for solution $g_{i}$ from CodeVerifier with the input $\text{TP}^{\prime}$
15:         if $r_{g_{i}}>l_{p}$ then
16:              Set $l_{p}=r_{g_{i}},l_{s}={g_{i}}$
17:         end if
18:     end for
19:     if $l_{p}>g_{p}$ then
20:         Set $g_{p}=l_{p},g_{s}=l_{s}$
21:         Add TP’ to TP
22:     end if
23:     if $g_{p}>\theta$ then
24:         return $g_{s}$ as the best solution
25:         Terminate optimal refinement iteration
26:     end if
27:end for
A.3 Exploration for Tests
@@
Table 8: The variation in the Pass@1 metric and the CodeVerifier’s overload is influenced by changes in the number of test cases $m$ during each exploration.
We randomly selected 60 tasks from the MBPP test data to observe the performance improvement and code execution overhead as the number of test cases $m$ increases during each exploration. Table 8 illustrates this trend. The base model we used is Kimi. We observed that increasing $m$ leads to improved Pass@1 performance and results in longer execution times due to the additional effort required for self-verification. We set $m$ for all experiments to balance time efficiency and performance.
A.4 Prompts for Exploration Agent
@@
A.5 Inputs for CodeVerifier
Figure 7: ThinkCoder CodeVerifier input example.
@@ A.5 Inputs for CodeVerifier
A.6 Example of Thorough Exploration
@@
A.7 Example of Optimal Refinement
@@
A.8 Example of Testing Pool